# Mapping surface height dynamics to subsurface flow physics in free-surface turbulent flow using a shallow recurrent decoder

### v1

@@ IV Results and discussion
In what follows, we present and discuss the results of reconstructing subsurface turbulent velocity fields from sparse measurements of the free surface only, by applying the SHRED algorithm to our turbulence data from the DNS and experimental cases. We first show and discuss the field reconstructions for shallow and deep horizontal planes, and compare these to the original and compressed fields. Secondly, we present a detailed analysis of the depth-dependent performance of SHRED, using the set of error metrics presented in Section III.4. Third, we demonstrate the capabilities of SHRED in temporal dynamics by looking at the RMS velocity time series of reconstructed fields. Finally, we show that the reconstructed fields yield reasonable turbulent power spectra for the relevant spatial scales, as compared to the compressed and ground-truth fields.
@@ IV.1 Reconstruction of surface elevation and velocity fields
To illustrate the performance of SHRED qualitatively, we present a side-by-side comparison of the surface elevation and a component of the horizontal velocity field: the original uncompressed data, their low-rank SVD approximations, and the SHRED reconstructions, for the S2 DNS case in Fig. 6 and the experimental case E2 in Fig. 7. The immediate eyeball observation is that the fields lose some sharpness in the SVD compression step, whereas the compressed and reconstructed fields are only distinguishable by eye upon careful inspection.
@@
From Fig. 6 one might get the impression that errors in the final reconstruction of the fields (right-hand column) occur during compression before training while the subsequent steps reproduce the compressed fields near-perfectly. To some extent this is correct. However, increasing the number of SVD modes in compression beyond the ranks chosen, does not necessarily improve the reconstruction. As we detail in Appendix A, training SHRED using the original uncompressed data set or with very little compression leads to considerable overfitting and far poorer results as well as higher computational expense.
@@
Inspecting the surface elevation for DNS case S2 in the top row of Fig. 6, we notice a sharp, dark, depression curve — referred to as a “scar” [1] — and brighter areas which are “boils” which signify upwelling of fluid to the surface. These features are successfully reconstructed, although the amplitudes can be seen to have been somewhat dampened. Scars also manifest in the surface velocity field of case S2, (middle row) as areas of fast flow. As with the surface elevation, the reconstruction looks visually indistinguishable from the low-rank approximation and the ground truth. Moreover, as demonstrated in the bottom row in Fig. 6, SHRED is capable of reconstructing the velocity field in the bulk flow, that is, at depths below the surface-influenced layer (depth for case S2; details on the surface-influenced layer in [2]).
@@
Similarly, for the experimental case E2, Fig. 7 shows visually good reconstruction performances across several planes. It is a considerably more challenging task to reconstruct planes in the experimental cases, due to noise and the much greater disparity of time and length scales at the far higher turbulent Reynolds numbers. However, we observe that even in this case, SHRED is capable of reconstructing from the time series of surface elevation in only three points, large and strong features such as scars at the surface elevation and medium-sized structures in the velocity fields as far down as cm () below the surface, although smaller features are somewhat blurred. More clearly than in the DNS case, velocity amplitudes are generally diminished in the reconstruction (right column) compared to the compressed training data (middle column).
@@
As an eyeball metric, Figs. 6 and 7 illustrate that SHRED is capable of reconstructing flow fields from sparse height measurements at the free surface in free-surface flows in both DNS simulations and experimental cases. SHRED’s performance is somewhat weaker in the experimental case, which is to be expected due to the presence of noise, greater range of turbulent scales, and significantly less data available for training compared to the long, continuous DNS datasets. In light of this we find it remarkable that SHRED reconstructs features of centimeter size cm beneath the surface, well outside the “blockage layer” where the free surface directly influences the velocity field [64, 2].
@@ IV.2 Depth-dependent performance
To quantify the performance of SHRED beyond simple comparisons, we use the five metrics discussed in section III.4 to evaluate different aspects of SHRED’s reconstruction of the horizontal velocity field . In addition to the averages taken in horizontal space and time we perform ensemble averaging over individual reconstructions for each flow case, where every ensemble case is constructed by a random distribution of the full data set into training, validation, and testing. The results are displayed with depth scaled by the integral length scale on the ordinate axis. Note that we adopt the convention from Aarnes et al. [2] of using an average measure for ‘horizontal’ grid plane depth for the DNS data, to allow for straightforward comparison of flow variables at a grid plane on the undulating grid without interpolation (details in [2]).
@@
Because the error metrics are averaged over time and ensembles, which could potentially conceal the direct performance error image-by-image, we include instantaneous profiles of Fig. 8, comparing reconstructions (dashed lines) to the uncompressed ground-truth values (solid lines) at three different arbitrarily chosen time instants. Note that for the two experimental cases, the measurements taken at different values of are not taken at the same time since different planes were measured separately. Generally, the reconstruction profiles match the ground truth profiles in shape but for a drop in magnitude, different for different depths and instants. There is a tendency for rapid changes in the measured profiles to be smoothed out in the reconstructed ones, particularly visible here for case S1 at instant 1. Such rapid changes are typically caused by intermittent turbulent events which are difficult for a neural network to learn, in particular for time-series-based networks such as the LSTM method used in SHRED. Strikingly, in all examples the accuracy of reconstruction deteriorates only very slightly with increasing depth.
@@
Shifting our attention to time-averaged results, the corresponding time-averaged quantities, , are plotted in panels (a) and (b) of Fig. 9 for DNS and experimental data, respectively, after averaging over an ensemble of 25 individual SHRED reconstructions each using a different set of three randomly chosen surface points as their sparse input. By normalizing the depth coordinate by the integral scale for each case, the profiles can be compared side-by-side. Again the solid lines represent the original data (i.e., ground truth) and color-matching dashed lines mark their reconstructed counterparts; the standard deviation of the ensemble is shown as shading (panel a) or error bars (panel b). As was observed at individual instants in Fig. 8, the trends of the profiles are retained by the reconstruction, but consistently offset to lower values, corresponding to a loss of energy in the reconstruction. This can only partly be explained by the compression which truncates the number of SVD modes; the offset is larger than expected by a rank reduction alone, as using a moderate truncation value for has a small impact on the large, energy-carrying scales of the flow (as seen in Fig. 3).
@@
The depth-variations of the Normalized Mean Squared Error (NMSE) of Eq. (16) are shown in Fig. 9(c). For the DNS cases, we notice a low MSE of about % close to the surface, and a steady increase up to around % for the lowermost planes. This is generally significantly lower than the subsurface reconstruction MSE of the CNN method of Xuan & Shen [18]. This might suggest that although both methods tend to result in dampened amplitudes of large-scale turbulence, SHRED performs better for this type of simulation data. The higher NMSE in the lower planes might be expected since there is hardly any correlation in a direct sense between the surface motion and the flow field at this depth.
@@
We observe that the PSD error metric (PSDE) follows similar trends as the NMSE. The errors are generally in the range of - in the large-to-intermediate length scales. The reconstructions of the DNS data show similar power spectral energy errors, although the high Re case S2 performs slightly better by this metric. The experimental case E1 shows lower errors than the DNS cases, which is also reflected in the PSD spectra of Fig. 11. This is, perhaps, unexpected because the data are generally noisier and harder to reconstruct, as highlighted by the other metrics. In contrast, the second experimental case, case E2, has the largest PSD error. We shall see below that, for all cases, the PSD error is always associated with a loss in turbulent kinetic energy in the reconstructions, at all lengthscales. Other turbulence-sensing reconstructions find similar losses in turbulent kinetic energy (see [18, 19]), although these are hard to compare directly with our results due to different flow cases and selection of error metrics.
@@
The SSIM results are presented in 9(e). The SSIM metric differs from the other error metrics that we use, as it targets structure, luminosity, and contrast, while the physical interpretation of the metric is not as obvious. The SSIM results reflect the qualitative observations made from Figs. 6 and 7, that there is an indisputable visual similarity between reconstructed fields and ground truth when the larger structures of the turbulent fields are considered. The DNS reconstructions yield SSIM values between and near the surface, falling to between and two integral length scales below the surface. For the experimental cases, peak values are found near the surface at around to , falling to around to in the deepest planes. While the latter might be considered a weak result in general image reconstruction schemes, not so in a turbulence context, considering that it is based on the time dynamics of three surface points only, after training. The SSIM likely emphasizes small-scale structures of the full-rank ground truth which are no longer present in the compressed training data (or, if they were, would be subject to overfitting; see Appendix A). This is particularly prominent for the SSIM results for experimental data. For the DNS data, the reconstruction SSIM of almost near the surface based on very sparse measurements, can be compared to super-resolution schemes for reconstructing full-state space from a coarse-grained and sparse field. The results of super-resolution reconstructions differ from case to case, but a case study on a turbulent LES found many methods to produce values of - [65]. In this regard, SHRED is close to matching this performance while also producing acceptable SSIM results for planes far from the surface.
@@
For the SHRED reconstructions, it is clear that the PSNR values, as seen in panel (f) of Fig. 9, are within acceptable values (above 20 dB) for all cases. We are not aware of other sufficiently similar reconstruction studies to which the values can be directly compared, yet one might note that super-resolution reconstructions of turbulent DNS data have found similar PSNR values [66]. The performance is matched, even with planes as deep as or further from the surface, with PSNR values of 22.5-25.5 dB. The DNS cases show higher values than the experimental ones, whereas the more turbulent ones show lower values. As expected, the PSNR generally decreases with depth, with the largest decrease occurring near the surface in all cases, where the flow changes rapidly due to surface viscous and blockage effects from the surface [64, 2].
@@ IV.3 Temporal analysis of planar Root-Mean-Square velocity
As an illustration of the SHRED reconstruction capabilities of temporal dynamics, we show the time series of the planar RMS values for a single SHRED reconstruction case, in Fig. 10. The planes chosen are relatively close to the surface, 1 cm depth for the experimental cases ( for E1 and for E2), and planes at depths for DNS cases S1 and S2. We observe a close correlation between the time series of the ground truth and the reconstruction. The normalized cross-correlation values with zero lag, are above for all cases except case E2, in which the correlation value is . The values of are generally lower for the reconstructed fields than the ground truth, indicating some loss of turbulent kinetic energy. Notably the error in does not noticeably increase when intermittent high-intensity turbulent events occur, corresponding to sudden peaks in the RMS velocity.
@@ IV.4 Spectral analysis
In this section, we investigate how the power density spectra of SHRED-reconstructed fields perform compared to ground truth and the low-rank SVD truncation. An example is shown in Fig. 11, for a SHRED case at depths of . The 1-D power density spectrum in the -direction, averaged over time and spatially in the direction, is calculated for each case. The purple dotted vertical line indicates the cutoff wavenumber for the low-rank truncation, i.e., the wavenumber at which the compressed spectrum deviates from ground truth by about %. (See section Section III.2 and Appendix A for details on selecting the rank truncation). One observes that generally, the spectra for the compressed field (dashed dotted red line) closely follow the spectra of the ground truth (solid dark line) up to the cutoff wavenumber. We observe that the spectra of the SHRED reconstructions generally follow the ground truth, but with a spectral energy loss evenly distributed over all frequencies.
@@ V Conclusions
In this study, we demonstrated a proof-of-concept application of the SHallow REcurrent Decoder (SHRED) for reconstructing subsurface turbulence fields from sparse surface measurements in free-surface flows. Using only three arbitrarily placed sensors measuring surface elevation as input, SHRED was able to infer the dominant large- and intermediate-scale structures below the surface across four different turbulent cases, two based on simulation data and two on experimental data.The differences between the data sets in terms of Reynolds number, sparsity, and noise demonstrate the flexibility of SHRED in handling different turbulence regimes.
@@
All code used in producing these results, are included and thoroughly presented in the following GitHub repository: https://github.com/krissmoe/SHRED-turbulence-sensing. The DNS data and supporting codes are available from [67]. Due to size constraints, raw PIV data is not hosted in the repo; download instructions are included in the GitHub README.
@@ Acknowledgements
DNS data was generously shared with us by Prof. Lian Shen and Dr. Anxing Xuan at the University of Minnesota. The work of JRA, KSM and SÅE was co-funded by the Research Council of Norway (iMOD, grant 325114) and the European Union (ERC CoG, WaTurSheD, grant 101045299). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The work of JNK was supported in part by the US National Science Foundation (NSF) AI Institute for Dynamical Systems (dynamicsai.org), grant 2112085. JNK further acknowledges support from the Air Force Office of Scientific Research (FA9550-24-1-0141).

### v2

@@ IV Results and discussion
In what follows, we present and discuss the results of reconstructing subsurface turbulent velocity fields from sparse measurements of the free surface only, by applying the SHRED algorithm to the turbulence data from the DNS [2] and experimental [4] datasets. We first show and discuss the field reconstructions for shallow and deep horizontal planes, and compare these to the original and compressed fields. Secondly, we present a detailed analysis of the depth-dependent performance of SHRED, using the set of error metrics presented in Section III.4. Third, we demonstrate the capabilities of SHRED in temporal dynamics by looking at the RMS velocity time series of reconstructed fields. Finally, we show that the reconstructed fields yield reasonable turbulent power spectra for the relevant spatial scales, as compared to the compressed and ground-truth fields.
@@ IV.1 Reconstruction of surface elevation and velocity fields
To illustrate the performance of SHRED qualitatively, we present a side-by-side comparison of the surface elevation and a component of the horizontal velocity field: the original uncompressed data, their low-rank SVD approximations, and the SHRED reconstructions, for the S2 DNS data in Fig. 6 and the experimental data E2 in Fig. 7. The immediate eyeball observation is that the fields lose some sharpness in the SVD compression step, whereas the compressed and reconstructed fields are only distinguishable by eye upon careful inspection.
@@
From Fig. 6 one might get the impression that errors in the final reconstruction of the fields (right-hand column) occur during compression before training while the subsequent steps reproduce the compressed fields near-perfectly. To some extent, this is correct. However, increasing the number of SVD modes in compression beyond the ranks chosen does not necessarily improve the reconstruction. As we detail in Appendix A, training SHRED using the original uncompressed dataset or with very little compression leads to considerable overfitting and far poorer results, as well as higher computational expense.
@@
Inspecting the surface elevation for DNS case S2 in the top row of Fig. 6, we notice a sharp, dark, depression curve — referred to as a “scar” [7] — and brighter areas which are “boils” which signify upwelling of fluid to the surface. These features are successfully reconstructed, although the amplitudes can be seen to have been somewhat dampened. Scars also manifest in the surface velocity field (middle row) as areas of fast flow. As with the surface elevation, the reconstruction looks visually indistinguishable from the low-rank approximation and the ground truth. Moreover, as demonstrated in the bottom row in Fig. 6, SHRED is capable of reconstructing the velocity field in the bulk flow, that is, at depths below the surface-influenced layer (depth for case S2; details on the surface-influenced layer in [1]).
@@
Similarly, for the experimental dataset E2, Fig. 7 shows visually good reconstruction performances across several planes. It is a considerably more challenging task to reconstruct planes in the experimental flows due to noise and the much greater disparity of time and length scales at the far higher turbulent Reynolds numbers. However, we observe that even in this case, SHRED is capable of reconstructing from the time series of surface elevation in only three points, large and strong features such as scars at the surface elevation and medium-sized structures in the velocity fields as far down as cm () below the surface, although smaller features are somewhat blurred. More clearly than in the DNS flow in Fig. 6, velocity amplitudes are generally diminished in the reconstruction (right column) compared to the compressed training data (middle column).
@@
As an eyeball metric, Figs. 6 and 7 illustrate that SHRED is capable of reconstructing flow fields from sparse height measurements at the free surface in free-surface turbulence data from DNS as well as state-of-the-art experiments. SHRED’s performance is somewhat weaker in the experimental case, which is to be expected due to the presence of noise, greater range of turbulent scales, and significantly less data available for training compared to the long, continuous DNS datasets. In light of this we find it remarkable that SHRED reconstructs features of centimeter size cm beneath the surface, well outside the “blockage layer” where the free surface directly influences the velocity field [50, 1].
@@ IV.2 Depth-dependent performance
To quantify the performance of SHRED beyond simple comparisons, we use the five metrics discussed in section III.4 to evaluate different aspects of SHRED’s reconstruction of the horizontal velocity field . In addition to the averages taken in horizontal space and time, we perform ensemble averaging over individual reconstructions for each flow case, where every time series in the ensemble is constructed by a random distribution of the full dataset into training, validation, and testing. The results are displayed with depth scaled by the integral length scale on the ordinate axis. Note that we adopt the convention from Aarnes et al. [1] of using an average measure for ‘horizontal’ grid plane depth for the DNS data, to allow for straightforward comparison of flow variables at a grid plane on the undulating grid without interpolation (details in [1]).
@@
Because the error metrics are averaged over time and ensembles, which could potentially conceal the direct performance error image-by-image, we include instantaneous profiles of Fig. 8, comparing reconstructions (dashed lines) to the uncompressed ground-truth values (solid lines) at three different arbitrarily chosen time instants. Note that for the two experimental flows, the measurements taken at different values of are not taken at the same time since different planes were measured separately. Generally, the reconstruction profiles match the ground truth profiles in shape but for a drop in magnitude, different for different depths and instants. There is a tendency for rapid changes in the measured profiles to be smoothed out in the reconstructed ones, particularly visible here for flow S1 at instant 1. Such rapid changes are typically caused by intermittent turbulent events which are difficult for a neural network to learn, in particular for time-series-based networks such as the LSTM method used in SHRED. Strikingly, in all examples the accuracy of reconstruction deteriorates only very slightly with increasing depth.
@@
Shifting our attention to time-averaged results, the corresponding time-averaged quantities, , are plotted in panels (a) and (b) of Fig. 9 for DNS and experimental data, respectively, after averaging over an ensemble of 25 individual SHRED reconstructions each using a different set of three randomly chosen surface points as their sparse input. By normalizing the depth coordinate by the integral scale for each flow, the profiles can be compared side-by-side. Again, the solid lines represent the original data (i.e., ground truth), and color-matching dashed lines mark their reconstructed counterparts; the standard deviation of the ensemble is shown as shading (panel a) or error bars (panel b). As was observed at individual instants in Fig. 8, the trends of the profiles are retained by the reconstruction, but consistently offset to lower values, corresponding to a loss of energy in the reconstruction. This can only partly be explained by the compression which truncates the number of SVD modes; the offset is larger than expected by a rank reduction alone, as using a moderate truncation value for has a small impact on the large, energy-carrying scales of the flow (as seen in Fig. 3).
@@
The depth-variations of the Normalized Mean Squared Error (NMSE) of Eq. (16) are shown in Fig. 9(c). For the DNS flows we notice a low MSE of about % close to the surface, and a steady increase up to around % for the lowermost planes. This is generally significantly lower than the subsurface reconstruction MSE of the CNN method of Xuan & Shen [64]. This might suggest that although both methods tend to result in dampened amplitudes of large-scale turbulence, SHRED performs better for this type of simulation data. The higher NMSE in the lower planes might be expected since there is hardly any correlation in a direct sense between the surface motion and the flow field at this depth.
@@
We observe that the PSD error metric (PSDE) follows similar trends as the NMSE. The errors are generally in the range of - in the large-to-intermediate length scales. The reconstructions of the DNS data show similar power spectral energy errors, although the high- data case S2 performs slightly better by this metric. The experimental flow E1 shows lower errors than the DNS cases, which is also reflected in the PSD spectra of Fig. 11. This is, perhaps, unexpected because the data are generally noisier and harder to reconstruct, as highlighted by the other metrics. In contrast, the second experimental flow, E2, has the largest PSD error. We shall see below that, for all datasets the PSD error is always associated with a loss in turbulent kinetic energy in the reconstructions, at all lengthscales. Other turbulence-sensing reconstructions find similar losses in turbulent kinetic energy (see [64, 13]), although these are hard to compare directly with our results due to different flow conditions and selection of error metrics.
@@
The SSIM results are presented in 9(e). The SSIM metric differs from the other error metrics that we use, as it targets structure, luminosity, and contrast, while the physical interpretation of the metric is not as obvious. The SSIM results reflect the qualitative observations made from Figs. 6 and 7, that there is an indisputable visual similarity between reconstructed fields and ground truth when the larger structures of the turbulent fields are considered. The DNS reconstructions yield SSIM values between and near the surface, falling to between and two integral length scales below the surface. For the experimental flows, peak values are found near the surface at around to , falling to around to in the deepest planes. While the latter might be considered a weak result in general image reconstruction schemes, not so in a turbulence context, considering that it is based on the time dynamics of three surface points only, after training. The SSIM likely emphasizes small-scale structures of the full-rank ground truth which are no longer present in the compressed training data (or, if they were, would be subject to overfitting; see Appendix A). This is particularly prominent for the SSIM results for experimental data. For the DNS data, the reconstruction SSIM of almost near the surface based on very sparse measurements, can be compared to super-resolution schemes for reconstructing full-state space from a coarse-grained and sparse field. The results of super-resolution reconstructions differ from case to case, but a study based on data from turbulent LES found many methods to produce values of - [11]. In this regard, SHRED is close to matching this performance while also producing acceptable SSIM results for planes far from the surface.
@@
For the SHRED reconstructions, it is clear that the PSNR values, as seen in panel (f) of Fig. 9, are within acceptable values (above 20 dB) for all datasets. We are not aware of other sufficiently similar reconstruction studies to which the values can be directly compared, yet one might note that super-resolution reconstructions of turbulent DNS data have found similar PSNR values [44]. The performance is matched, even with planes as deep as or further from the surface, with PSNR values of 22.5-25.5 dB. The DNS datasets yield higher values than the experimental ones, whereas the more turbulent ones show lower values. As expected, the PSNR value generally decreases with depth, with the largest decrease occurring near the surface in all cases, where the flow changes rapidly due to surface viscous and blockage effects from the surface [50, 1].
@@ IV.3 Temporal analysis of planar Root-Mean-Square velocity
As an illustration of the SHRED reconstruction capabilities of temporal dynamics, we show the time series of the planar RMS values for a single SHRED reconstruction case, in Fig. 10. The planes chosen are relatively close to the surface, 1 cm depth for the experimental flows ( for E1 and for E2), and planes at depths for DNS flow cases S1 and S2. We observe a close correlation between the time series of the ground truth and the reconstruction. The normalized cross-correlation values with zero lag, are above for all datasets except E2, in which the correlation value is . The values of are generally lower for the reconstructed fields than the ground truth, indicating some loss of turbulent kinetic energy. Notably the error in does not noticeably increase when intermittent high-intensity turbulent events occur, corresponding to sudden peaks in the RMS velocity.
@@ IV.4 Spectral analysis
In this section, we investigate how the power density spectra of SHRED-reconstructed fields perform compared to ground truth and the low-rank SVD truncation. An example is shown in Fig. 11, for a SHRED case at depths of . The 1-D power density spectrum in the -direction, averaged over time and spatially in the direction, is calculated in each case. The purple dotted vertical line indicates the cutoff wavenumber for the low-rank truncation, i.e., the wavenumber at which the compressed spectrum deviates from ground truth by about %. (See section Section III.2 and Appendix A for details on selecting the rank truncation). One observes that generally, the spectra for the compressed field (dashed dotted red line) closely follow the spectra of the ground truth (solid dark line) up to the cutoff wavenumber. We observe that the spectra of the SHRED reconstructions generally follow the ground truth, but with a spectral energy loss evenly distributed over all frequencies.
@@ V Conclusions
In this study, we demonstrated a proof-of-concept application of the SHallow REcurrent Decoder (SHRED) for reconstructing subsurface turbulence fields from sparse surface measurements in free-surface flows. Using only three arbitrarily placed sensors measuring surface elevation as input, SHRED was able to infer the dominant large- and intermediate-scale structures below the surface across four different datasets of turbulent flow beneath a free surface together with the instantaneous elevation of the surface itself, two based on simulation data and two on experimental data recently published in [1] and [3]. The differences between the datasets in terms of Reynolds number, sparsity, and noise demonstrate the flexibility of SHRED in handling different turbulence regimes.
@@
All code used in producing these results, are included and thoroughly presented in the following GitHub repository: https://github.com/krissmoe/SHRED-turbulence-sensing. The DNS data and supporting codes are available from [2], the experimental data are available from [4].
@@ Acknowledgements
The experimental datasets are from an experiment performed by Ali Semati, Dr. Amélie Ferran and Dr. Yi Hui Tee under the guidance of Prof. R. Jason Hearst; beyond the experimental data itself we have benefited from discussions with these and input on the manuscript. DNS data was generously shared with us by Prof. Lian Shen and Dr. Anxing Xuan at the University of Minnesota. We thank Omer M. Babiker for assistance with data handling and preparation, and many discussions. The work of KSM, JRAa and SÅE was co-funded by the Research Council of Norway (iMOD, grant 325114) and the European Union (ERC CoG, WaTurSheD, grant 101045299). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The work of JNK was supported in part by the US National Science Foundation (NSF) AI Institute for Dynamical Systems (dynamicsai.org), grant 2112085. JNK further acknowledges support from the Air Force Office of Scientific Research (FA9550-24-1-0141).
Author contributions
KSM and JRAa performed the data analysis with SHRED and wrote the first draft of the majority of the manuscript. JNK made the first implementation, co-developed the concept and contributed writing and supervision. SÅE contributed to the concept and discussions, supervision. All authors contributed significantly to the writing of the final manuscript.